OpenStudy (jacobbirddog):

Please help, I need help on these last few questions I will fan and medal you.

OpenStudy (jacobbirddog):

Find the lengths UQ, VR, and TP on triangle SWN if segment WN = 32 ft.

OpenStudy (jacobbirddog):

please help!

OpenStudy (taylor0402):

Corresponding parts of the triangles are in proportion. By examination we see that the sides of the larger triangle are divided into four equal segments. So, for example, ST / SW = 3/4. ST / SW = TP / NW 3/4 = TP/26 TP = 26*3/4 = 39/2 = 19.5 ft. In similar fashion we have: SU / SW = UQ / NW 2/4 = TP/26 TP = 2*26/4 = 13 ft. SV/SW = VR/NW 1/4 = VR/26 VR = 26/4 = 13/2 = 6.5 ft. Use the examples i have provided above to solve the answer to your problem

OpenStudy (taylor0402):

Hope this helps 0_o

OpenStudy (jacobbirddog):

yes it does!